OpenStudy (anonymous):

If x=3 and y=4, find the value of 2x+y-1

OpenStudy (jamesj):

For real? Substitute in the values 2x + y + 1 = 2 times 3 + 4 + 1 = 6 + 4 + 1 = ...

OpenStudy (anonymous):

ohhh,,, i was thinkin something totally else i know it now